OpenStudy (hba):

1 * 1 * 1 * 0 - 0 * 2 +2 = ?

OpenStudy (anonymous):

PEMDAS?

OpenStudy (anonymous):

0- 0 + 2?

OpenStudy (hba):

how do we do it its very diff

OpenStudy (anonymous):

PEMDAS is when you do parentheses exponents multiplication division addition and subtraction so we start with the multiplication 1*1*1*1*0 = 0 0 *2 = 0 So we have 0-0 + 2 0-0=0 + 2 = 2

OpenStudy (hba):

bye bye ppl

OpenStudy (anonymous):

?

OpenStudy (hba):

i really didnt get it

OpenStudy (hba):

@Compassionate

OpenStudy (hba):

i know bodmas

OpenStudy (anonymous):

Oh well you should say that xD PEMDAS is the order of operations You do problems in this order. You always start with parentheses then move to exponents the multiplication then division then additino and lastly subtraction. If any of these are not present in a problem you skip it. This problem only has multiplication subtraction and addition 1 times itself is always 1 anything times 0 is 0

OpenStudy (hba):

i am very keen to learn but please explain me

OpenStudy (hba):

:D

OpenStudy (hba):

@rebeccaskell94 i didnt know that really ur so intelligent

OpenStudy (anonymous):

lol thanks. It's no big deal xD let me know if you need help with anything else :)

OpenStudy (compassionate):

PEMDAS: P arenthesis E exponents M multiplication D division A addition S subtraction 1 * 1 * 1 * 0 - 0 * 2 +2 Multiplication first: 1 x 1 x 1 x 0 = 0 20 now we have 0 - 0 x 2 + 2 Multiply 0 x 2 = 0 Now we have 0 - 0 + 2 Now do the addition: 0 + 2 = 2 Now 0 - 2 = 2 ANS: 2

OpenStudy (hba):

[; \int ;] what is this

OpenStudy (hba):

@rebeccaskell94

OpenStudy (anonymous):

That is latex for chat. You'll have to ask the people who use it haha

OpenStudy (hba):

:D

OpenStudy (compassionate):

ANS = - 2

OpenStudy (compassionate):

Not +2

OpenStudy (hba):

comp ur right i was testing @rebeccaskell94

OpenStudy (compassionate):

Did I pass? :)

OpenStudy (hba):

yes u did n u git ur medal for it

OpenStudy (hba):

@rebeccaskell94 please make ur base strong